# Tug-of-war between idioms’ figurative and literal interpretations in LLMs

### v4

@@ Abstract
Idioms present a unique challenge for language models due to their non-compositional figurative interpretations, which often strongly diverge from the idiom’s literal interpretation. In this paper, we employ causal tracing to systematically analyze how pretrained causal transformers deal with this ambiguity. We localize three mechanisms: (i) Early sublayers and specific attention heads retrieve an idiom’s figurative interpretation, while suppressing its literal interpretation. (ii) When disambiguating context precedes the idiom, the model leverages it from the earliest layer and later layers refine the interpretation if the context conflicts with the retrieved interpretation. (iii) Then, selective, competing pathways carry both interpretations: an intermediate pathway prioritizes the figurative interpretation and a parallel direct route favors the literal interpretation, ensuring that both readings remain available. Our findings provide mechanistic evidence for idiom comprehension in autoregressive transformers [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License..
@@ 2 Idiom interpretation task
We formalize a next-token prediction task and construct a dataset. Idioms are embedded into a sentence template designed to reveal their interpretation through continuation ‘X (would) IDIOM because X was so/too/a/the’ [^2] Here, X is instantiated with pronouns and morphologically adapted; X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were.. This template preserves idiom ambiguity while eliciting causal completions (e.g., for ‘kick the bucket’, figurative completions include ‘old/sick’, while literal completions include ‘angry/mad’.) Since there is no single ground truth continuation, we automatically construct token sets consistent with either the figurative or literal interpretation. The tokens in each set are mutually exclusive and chosen for their high likelihood of association with the respective interpretation. Below, we outline the dataset construction process, and see Figure 9:
@@
2. Sentence Generation: We generate a literal paraphrase and a figurative paraphrase for each idiom using the Llama3.3-70B-instruct model Grattafiori et al. (2024). The sentences below are three examples of the generated ambiguous sentence ($s_{a}$), figurative paraphrase ($s_{f}$), and literal paraphrase ($s_{l}$).
@@ 3.1 Information flow in transformers
Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathbf{x}_{i}^{0}=\mathrm{PE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{PE}$ denotes a positional encoding function that injects position $i$ into the token embedding Vaswani et al. (2017), and $d$ is the model’s hidden dimension. From layer 0, these vectors $\mathrm{x}_{i}^{0}$ are carried forward and accumulated over the subsequent $L$ layers. At each layer $\ell$, the hidden state for token $i$ is updated from its previous value $\mathrm{x}_{i}^{\ell-1}$ by adding the multi-head self-attention (MHSA) output $a_{i}^{\ell}$ and MLP output $m_{i}^{\ell}$:
@@ 4.1 Probing sublayers through knockout
To locate the mechanisms for retrieving figurative interpretations from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$. By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change. To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to the MLP and MHSA in each block.
@@
Results Figure 2(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out. We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2). At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero. And these shifts are significantly different in 0-3 layers ($p<0.05$). This indicates that early layer MLPs are causally important for reading out the figurative interpretation while potentially suppressing the literal one. For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, where the corresponding patching effect converges to zero.
@@ 4.2 Identifying attention heads specific to retrieving figurative interpretations
Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative interpretation. We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$.
@@
Figure 3: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span. Idiomatic heads: Heads those are crucial for retrieving the figurative interpretation of idiom; $\Delta F(s_{a})$ and $\Delta L(s_{a})$.
To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$. Figure 3 shows that there is a subset of attention heads that simultaneously exhibits large $\Delta F(s_{a})$ (Figure 3(a)) and $\Delta L(s_{a})$ (Figure 3(b)). To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples for each head’s contribution to the interpretation. We then rank all heads by, on the one hand, $\Delta F(s_{a})$, and, on the other hand, $\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders (selected attention heads listed in Appendix Table 1).
@@ 4.3 Causal role of MLP sublayers and idiomatic heads in figurative reading
To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch the activations on $s_{a}$ (at idiom span) into the corresponding components when processing $s_{l}$. As a control, we patch randomly chosen MLP sublayers and heads.
@@ 5.2 Analyzing layer-wise competing interpretation flow through activation patching
We interpret this as showing that a figurative injection doesn’t lead to suppression of the literal interpretation (an asymmetry between increases and decreases). By contrast, a literal injection must first suppress the figurative interpretation in that subspace to have an effect, producing a more symmetric push–pull tradeoff. This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant route for the figurative interpretation ( 3-1 figurative path).
@@
To probe this, we replace the idiom sentence’s idiom token span activation with the activation from a paraphrased sentence, and re-patch the ‘because’ token with the idiom sentence’s original activation (as shown in Figures 5(c), 5(d)). This preserves the intermediate pathway and isolates the contribution of the direct route. Figure 6(c) shows that removing direct path’s literal interpretation produces a large drop in $\Delta L$ (min: 0.25-0.25) and rise in $\Delta F$ (max: $0.20$). Conversely, removing the direct path’s figurative interpretation yields only a modest drop in $\Delta F$ (min: 0.10-0.10) and a small increase in $\Delta L$ (max: $0.10$). These results indicate that the direct path predominantly conveys literal information. The pattern is consistent across different models (Appendix E).
@@ 6 Localizing idiom disambiguation in context
Figure 7: Sublayer-wise interpretation shift $\Delta I$ after knockout activations at idiom span, for contexts FCFC: Figurative Context; LCLC: Literal Context.
ResultsFigure 7(a) shows that for idioms in a figurative context, knocking out the MLP in the early layers (0–3) results in a significant drop in $\Delta F$ (min: 0.62-0.62) and increase in $\Delta L$ (max: $0.66$), consistent with the pattern observed for idioms without context (Figure 2(a)). This indicates that these layers are critical for retrieving the figurative interpretation independently of context. This result is consistent across different models (Appendix F). For idioms in a literal context (7(b)), ablating the MLP layer 1 component leads to a significant drop in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.09$), indicating that the figurative interpretation is still being retrieved. The effect persists but is small compared with the figurative context condition, where early layers at the idiom token appear to encode context that suppresses figurative retrieval while biasing toward the literal interpretation.
@@
In sum, the results point to a two stage context disambiguation: (a) early layers retrieve figurative evidence while beginning to encode context, followed by (b) mid-layers, MHSA-driven contextual disambiguation that resolves conflicts in favor of the context-appropriate interpretation.
6.2 Probing MHSA-mediated contextual disambiguation through KVQ patching
To examine in detail how MHSA mediates contextual disambiguation in early-mid layers, we perform KVQ patching, in which MHSA’s key–value (KV) activations are replaced while selectively preserving idiom tokens’ query (Q) activations across layers, thereby steering the model’s attention toward alternative contexts. That is, we redirect idiom token span’s attention from a figurative-context to the literal context, and vice versa. If the model’s final prediction changes from figurative to literal (or vice versa) after patching, it demonstrates that the MHSA mechanism in those specific layers is responsible for using the context to steer the disambiguation process.
[…]
@@ 7 Conclusion
We present a multi-stage mechanism of idiom comprehension in causal transformers, by identifying distinct components and flows for figurative and literal interpretations. Causal interventions reveal (i) early retrieval of the idiomatic interpretation, (ii) immediate use of preceding context, with later refinement when it conflicts with the retrieved interpretation, (iii) a selective pathway that carries the figurative interpretation through intermediate layers, and a bypass route that favorably delivers the literal interpretation directly to the output. To sum up, these findings link representational dynamics to contextual disambiguation, yielding a concise mechanistic interpretation of idiom processing.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Starting Grant “Individualized Interaction in Discourse”, grant agreement No. 948878). Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102.

### v5

@@ Abstract
Idioms present a unique challenge for language models due to their non-compositional figurative interpretations, which often strongly diverge from the idiom’s literal interpretation. In this paper, we employ causal tracing to systematically analyze how pretrained causal transformers deal with this ambiguity. We localize three mechanisms: (i) Early sublayers and specific attention heads retrieve an idiom’s figurative interpretation, while suppressing its literal interpretation. (ii) When disambiguating context precedes the idiom, the model leverages it from the earliest layer and later layers refine the interpretation if the context conflicts with the retrieved interpretation. (iii) Then, selective, competing pathways carry both interpretations: an intermediate pathway prioritizes the figurative interpretation and a parallel direct route favors the literal interpretation, ensuring that both readings remain available. Our findings provide mechanistic evidence for idiom comprehension in autoregressive transformers [^1] Code and dataset are available at https://github.com/sori424/idiom_processing.
@@ 2 Idiom interpretation task
We formalize a next-token prediction task and construct a dataset. Idioms are embedded into a sentence template designed to reveal their interpretation through continuation ‘X (would) IDIOM because X was so/too/a/the’ [^2] Here, X is instantiated with pronouns and morphologically adapted; X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were.. This template preserves idiom ambiguity while eliciting causal completions (e.g., for ‘kick the bucket’, figurative completions include ‘old/sick’, while literal completions include ‘angry/mad’.) Since there is no single ground truth continuation, we automatically construct token sets consistent with either the figurative or literal interpretation. The tokens in each set are mutually exclusive and chosen for their high likelihood of association with the respective interpretation. Below, we outline the dataset construction process, and see Figure 9.
@@
2. Sentence Generation: We generate a literal paraphrase and a figurative paraphrase for each idiom using the Llama3.3-70B-instruct model Grattafiori et al. (2024). The sentences below are three examples of the generated ambiguous sentence ($s_{a}$), figurative paraphrase ($s_{f}$), and literal paraphrase ($s_{l}$):
@@ 3.1 Information flow in transformers
Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{N}]$ over a vocabulary $V$, each token is embedded as $\mathbf{x}_{i}^{0}=\mathrm{PE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{PE}$ denotes a positional encoding function that injects position $i$ into the token embedding Vaswani et al. (2017), and $d$ is the model’s hidden dimension. From layer 0, these vectors $\mathrm{x}_{i}^{0}$ are carried forward and accumulated over the subsequent $L$ layers. At each layer $\ell$, the hidden state for token $i$ is updated from its previous value $\mathrm{x}_{i}^{\ell-1}$ by adding the multi-head self-attention (MHSA) output $a_{i}^{\ell}$ and MLP output $m_{i}^{\ell}$:
@@ 4.1 Probing sublayers through knockout
To locate the mechanisms for retrieving figurative interpretations, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$. By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change. To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to a MLP and MHSA in each block.
@@
Figure 2: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for sentences $s\in\{s_{a},s_{f},s_{l}\}$. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals across idiom sentences. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
Results Figure 2(a) shows by how much the model’s interpretation of the idiom changes when the MLP sublayer is knocked out. We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2). At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero. And these shifts are significantly different in 0-3 layers ($p<0.05$). This indicates that early layer MLPs are causally important for reading out the figurative interpretation while potentially suppressing the literal one. For later layers ($\geq$ 4), the knockout does not affect the model’s interpretation, where the corresponding patching effect converges to zero.
@@ 4.2 Identifying attention heads specific to retrieving figurative interpretations
Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative interpretation. We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$. To operationalize this, we define idiomatic heads ($\mathcal{H}_{\mathrm{idiom}}$) as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$.
@@
Figure 3: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span. Idiomatic heads: Heads those are crucial for retrieving the figurative interpretation of idiom; $\Delta F(s_{a})$ and $\Delta L(s_{a})$, simultaneously.
We identify attention heads whose removal reliably decreases figurative probability ($\Delta F(s_{a})$), while increasing literal probability ($\Delta L(s_{a})$) for the idiom sentences. To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples for each head’s contribution to the interpretation. We then rank all heads by, on the one hand, $\Delta F(s_{a})$, and, on the other hand, $\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders (selected attention heads listed in Appendix Table 1). Figure 3 shows that there is a subset of attention heads that simultaneously exhibits large $\Delta F(s_{a})$ (Figure 3(a)) and $\Delta L(s_{a})$ (Figure 3(b)).
@@ 4.3 Causal role of MLP sublayers and idiomatic heads in figurative reading
To show that the components that we found (i.e., early MLP sublayers and $\mathcal{H}_{\mathrm{idiom}}$) are causally sufficient for figurative reading, we patch the activations on $s_{a}$ at idiom span tokens into the corresponding components when processing $s_{l}$: patching their activations from the idiom sentence into the literal paraphrase increases $\Delta F(s_{l}\hookleftarrow s_{a})$ and decreases $\Delta L(s_{l}\hookleftarrow s_{a})$, while patching random components has no effect.
@@ 5.2 Analyzing layer-wise competing interpretation flow through activation patching
The asymmetry between increases and decreases show that a figurative injection doesn’t lead to suppression of the literal interpretation. By contrast, a literal injection must first suppress the figurative interpretation in that subspace to have an effect, producing a more symmetric push–pull tradeoff. This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant route for the figurative interpretation ( 3-1 figurative path).
@@
To probe this, we replace the idiom sentence’s idiom token span activation with the activation from a paraphrased sentence, and re-patch the ‘because’ token with the idiom sentence’s original activation (as shown in Figures 5(c), 5(d)). This preserves the intermediate pathway and isolates the contribution of the direct route. Figure 6(c) shows that removing direct path’s literal interpretation produces a large drop in $\Delta L$ (min: 0.25-0.25) and rise in $\Delta F$ (max: $0.20$). Conversely, removing the direct path’s figurative interpretation yields only a modest drop in $\Delta F$ (min: 0.10-0.10) and a small increase in $\Delta L$ (max: $0.10$). These results indicate that the direct path predominantly conveys literal information. This pattern is consistent across different models (Appendix E).
@@ 6 Localizing idiom disambiguation in context
Figure 7: Sublayer-wise interpretation shift $\Delta I$ after knockout activations at idiom span with contexts. The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
ResultsFigure 7(a) shows that for idioms in a figurative context, knocking out the MLP in the early layers (0–3) results in a significant drop in $\Delta F$ (min: 0.62-0.62) and increase in $\Delta L$ (max: $0.66$), consistent with the pattern observed for idioms without context (Figure 2(a)). This indicates that these layers are critical for retrieving the figurative interpretation independently of context. This result is consistent across different models (Appendix F). For idioms in a literal context (7(b)), ablating the MLP layer 1 leads to a significant drop in $\Delta F$ (min: 0.33-0.33) and an increase in $\Delta L$ (max: $0.09$), indicating that the figurative interpretation is still being retrieved. The effect persists but is small compared with the figurative context condition, where early layers at the idiom token appear to encode context that suppresses figurative retrieval while biasing toward the literal interpretation.
@@
In sum, the results point to a two stage context disambiguation: (1) early layers retrieve figurative evidence while beginning to encode context, followed by (2) mid-layers, MHSA-driven contextual disambiguation that resolves conflicts in favor of the context-aligned interpretation.
6.2 Probing MHSA contextual disambiguation with Query-preserved KV patching
To examine in detail how MHSA mediates contextual disambiguation in early-mid layers, we perform Query-preserved Key-Value patching, in which MHSA’s Key–Value (KV) activations are swapped while selectively preserving idiom tokens’ query (Q) activations across layers, thereby redirecting the model’s attention toward alternative contexts. That is, we switch idiom token span’s attention from a figurative-context to the literal context, and vice versa. If the model’s final prediction changes from figurative to literal (or vice versa) after patching, it demonstrates that the MHSA mechanism in those specific layers is responsible for using the context in the disambiguation process.
(a) FC $s_{a} \hookleftarrow$ LC $s_{a}$
at idiom span
(b) LC $s_{a} \hookleftarrow$ FC $s_{a}$
at idiom span
Figure 8: Layer-wise interpretation shifts $\Delta I$ under Q-preserved KV patching. Y-axis: Mean $\Delta L(s_{a})$ and $\Delta F(s_{a})$ across idiom sentences. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ exceeds the average difference across layers (paired $t$-test, $p<0.05$).
@@ 7 Conclusion
We present a multi-stage mechanism of idiom comprehension in causal transformers, by identifying distinct components and flows for figurative and literal interpretations. Causal interventions reveal (i) early retrieval of the idiomatic interpretation, (ii) immediate use of preceding context, with later refinement when it conflicts with the retrieved interpretation, (iii) a selective pathway that carries the figurative interpretation through intermediate layers, and a bypass route that favorably delivers the literal interpretation directly to the output.
Considering idioms as multiword expressions, our results suggest that early layers (especially MLPs) perform a “detokenization” step Gurnee et al. (2023); Elhage et al. (2021) that compresses the idiom into a unified figurative representation, while later pathways can “retokenize” it into a competing literal interpretation. To sum up, these findings link representational dynamics to contextual disambiguation, yielding a concise mechanistic interpretation of idiom processing.
@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Starting Grant “Individualized Interaction in Discourse”, grant agreement No. 948878). Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. We gratefully acknowledge the stimulating research environment of the GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project number 471607914.